# Robust Superradiance and Spontaneous Spin Ordering in Disordered Waveguide Quantum Electrodynamics

### v2

@@ VII Conclusion and Outlook
Looking forward, several promising directions emerge. While our current model involves two decay channels arising from the bidirectional waveguide, extensions to systems with multiple decay channels may reveal new forms of collective phenomena. Additionally, studying superradiance in engineered environments—such as squeezed reservoirs, structured bath, or chiral waveguides—may uncover richer phenomena of coherence and self-organization. The analytical and numerical frameworks introduced here are broadly adaptable for future studies of disorder, coherence, and collectivity in many-body quantum optics. Beyond its fundamental interest, the predicted robustness of these collective quantum effects against various forms of local disorders is also of significant practical importance for potential applications that rely on collective radiation effects. This includes the efficient generation of metrologically useful states [46, 47, 48, 49], the development of ultra-narrow superradiant lasers [43, 44, 45], as well as recent proposals for superradiant quantum batteries [74]. Our findings suggest that such applications exhibit significantly greater resilience to imperfections than commonly assumed.
@@ Acknowledgements.
We thank Iman Marvian for a helpful conversation. We acknowledge support by the Deutsche Forschungsgemeinschaft (German Research Foundation)–522216022 and by the ERC project ‘ConsQuanDyn’ (No. 851161). D. M. acknowledges support from the Novo Nordisk Foundation under Grants No. NNF22OC0071934 and No. NNF20OC0059939. This research is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.
@@ Appendix A Finite-Size Scaling of the Decay Rates
where $\hat{R}$ is a Hermitian matrix defined by the jump operators in Eq. (3), $\Gamma_{k}$ are its eigenvalues, and $\Omega_{k}=\sum_{j}U_{kj}\sigma_{j}^{-}$ are the corresponding decay modes after diagonalization of $\hat{R}$ with a unitary transformation $U$. While the eigenvalues are only directly related to the decay rates of the first few emissions, an upper bound of the maximal decay rate can be written out as (see Appendix E)
@@ Appendix C Numerical Benchmarking of Different Methods
Appendix D Gaussian-Distributed Positions
@@
Appendix E A Loose Upper Bound
@@
Appendix F Neglecting the Hamiltonian
@@
Appendix G Spontaneous Spin Ordering in a Regular Lattice
@@
Appendix H Time Evolution of Left- and Right-Emission Rates
@@
As a complement to Fig. 9, Fig. 16(a) shows the time evolution of the left-right emission asymmetry $(\mathcal{R}_{\text{R}}-\mathcal{R}_{\text{L}})$ for individual QSDMF trajectories for $\Theta=2\pi$. It can be seen that most trajectories exhibit an asymmetry in the emission rate with a preference for either one or the other side. However, this asymmetry is not completely fixed during the decay and some of the trajectories change their preferred direction of emission during the decay process. As shown in Fig. 16(b), relative strength of the asymmetry peaks at $t_{\star}$ and increases with disorder strength.
Appendix I Boundary Effects
@@
Figure S3: (a) Probability distribution of relative propagating phases $\Delta\xi=(\xi_{i}-\xi_{j})\bmod 2\pi$, which is a wrapped triangular distribution. (b) Probability distribution of relative spin angles $\Delta\phi=(\phi_{i}-\phi_{j})\bmod 2\pi$, for $N=400$ and $10^{3}$ trajectories. When $\Theta=0$, the system is permutation-invariant in QSDMF trajectories, and all spins align along the same direction, resulting in $p(\Delta\phi)=\delta(0)$ as expected. As $\Theta$ increases, the relative propagating phase $\Delta\xi$ becomes more uniformly distributed in $(-\pi,\pi)$. Accordingly, $p(\Delta\phi)$ also becomes more evenly distributed, meaning the global spin alignment of the atoms weakens. Instead, a new spin ordering emerges, with spins aligning according to their position-dependent phases as revealed by the joint probability distribution $p(\Delta\xi,\Delta\phi)$ shown in Sec. V.2. Additionally, we can observe that, under strong disorder, there is a weak tendency for alignment and anti-alignment in the distribution $p(\Delta\phi)$. This can be attributed to the fact that the joint distribution $p(\Delta\xi,\Delta\phi)$, shown in Fig. 8, has less spreading around $\Delta\xi\approx 0$ and $\pi$. For similar reasons as in Appendix G, the ordering around $\Delta\phi\approx\Delta\xi\approx\pm\pi/2$ is weaker than ordering in other directions.
@@
K Reduction of the SDE for Efficient Treatment of Cavity Dynamics
@@
L Irrelevance of Detuning in DTWA

### v3

@@ VII.3 Outlook
Looking forward, several promising directions emerge. While our current model involves two decay channels arising from the bidirectional waveguide, extensions to systems with multiple decay channels may reveal new forms of collective phenomena. Though the synchronization process can be captured quantitatively by our semiclassical EOMs, a transparent and qualitative theory describing the dynamical formation of spin ordering would also be interesting for future research. Additionally, studying superradiance in engineered environments—such as squeezed reservoirs, structured bath, or chiral waveguides—may uncover richer phenomena of coherence and self-organization. The analytical and numerical frameworks introduced here are broadly adaptable for future studies of disorder, coherence, and collectivity in many-body quantum optics. Beyond its fundamental interest, the predicted robustness of these collective quantum effects against various forms of local disorders is also of significant practical importance for potential applications that rely on collective radiation effects. This includes the efficient generation of metrologically useful states [53, 54, 55, 56, 57], the development of ultra-narrow superradiant lasers [50, 51, 52], as well as recent proposals for superradiant quantum batteries [83]. Our findings suggest that such applications exhibit significantly greater resilience to imperfections than commonly assumed. More broadly, the existence of cooperative effects that remain robust even under strong disorder, such as the spin ordering identified here, challenges the conventional intuition that imperfections destroy nontrivial collective quantum behavior in realistic platforms.
@@ Acknowledgements.
We thank Constanze Bach, Iman Marvian, Arno Rauschenbeutel, and Philipp Schneeweiss for helpful discussions. We acknowledge support by the Deutsche Forschungsgemeinschaft (German Research Foundation)–522216022 and by the ERC project ‘ConsQuanDyn’ (No. 851161). D. M. acknowledges support from the Novo Nordisk Foundation under Grants No. NNF22OC0071934 and No. NNF20OC0059939. This research is part of the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.
@@ Appendix A Finite-Size Scaling of the Decay Rates
where $\hat{R}$ is a Hermitian matrix defined by the jump operators in Eq. (3), $\Gamma_{k}$ are its eigenvalues, and $\Omega_{k}=\sum_{j}U_{kj}\sigma_{j}^{-}$ are the corresponding decay modes after diagonalization of $\hat{R}$ with a unitary transformation $U$. While the eigenvalues are only directly related to the decay rates of the first few emissions, an upper bound of the maximal decay rate can be written out as (see Appendix F)
@@ Appendix C Numerical Benchmarking of Different Methods
Appendix D Vanishing Mutual Information in Disordered Waveguide QED
To further substantiate the validity of QSDMF in the large-$N$ limit, we verify the finding of Ref. [43]—namely that the mutual information along QSD trajectories vanishes with increasing system size—also in the present, strongly disordered setting. This property justifies the validity of the product-state approximation at the trajectory level.
In Fig. 13, we perform MPS simulations of the QSD trajectories for the representative case $\Theta=2\pi$. For $N=50$, the trajectory-averaged bipartite entanglement entropy $\tilde{S}_{N/2}$ requires a bond dimension $d\sim 20$ for convergence. In contrast, local observables, such as the decay rate $\mathcal{R}$, are already reproduced by the MF approximation with negligible error. The reason lies in the fact that the trajectory-averaged mutual information between a pair of spins $\tilde{I}_{j,j^{\prime}}$ becomes vanishingly small with increasing $N$. In Fig. 13(c) and (d), we show $\tilde{I}_{j,j^{\prime}}$ for the central pair of TLAs at the superradiant peak time $t=t_{\star}$ and at $t=1/\gamma$ in the subradiant regime, respectively. In both cases, although $\tilde{S}_{N/2}$ grows with $N$, $\tilde{I}_{N/2,N/2+1}$ remains far below unity and continues to decrease. Other spin pairs behave similarly. These results, consistent with Ref. [43], show that QSDMF remains highly accurate for local observables at large $N$.
Figure 13: Vanishing mutual information. MPS simulations are performed for QSD trajectories with increasing bond dimension $d$ at $\Theta=2\pi$. Converged results are already obtained for a bond dimension of $d=20$. (a) Time evolution of the decay rate $\mathcal{R}/N$ for $N=50$. The vertical line marks the superradiant peak time $t_{\star}$, and the pink shaded region denotes the subradiant regime. The difference between the QSDMF result ($d=1$) and the numerically exact result is negligible, thereby verifying the accuracy of QSDMF for $N=50$. (b) Time evolution of the trajectory-averaged bipartite entanglement entropy $\tilde{S}_{N/2}$ and the trajectory-averaged mutual information between the two central TLAs, $\tilde{I}_{N/2,N/2+1}$, for $N=50$. (c) and (d) Scaling of $\tilde{S}_{N/2}$ and $\tilde{I}_{N/2,N/2+1}$ with NN at the superradiant peak time $t_{\star}$ and at $\gamma t=1$ in the subradiant regime, respectively. In both cases, $d=20$ is used, and the mutual information between the two TLAs decreases as NN increases. All simulations have been performed with a time step $\gamma\delta t=2\times 10^{-4}$ and using $\mathcal{N}=10^{3}$ trajectories.
Appendix E Gaussian-Distributed Positions
@@
Appendix F A Loose Upper Bound
@@
Appendix G Neglecting the Hamiltonian
@@
Appendix H Spontaneous Spin Ordering in a Regular Lattice
@@
Appendix I Time Evolution of Left- and Right-Emission Rates
@@
As a complement to Fig. 9, Fig. 17(a) shows the time evolution of the left-right emission asymmetry $(\mathcal{R}_{\text{R}}-\mathcal{R}_{\text{L}})$ for individual QSDMF trajectories for $\Theta=2\pi$. It can be seen that most trajectories exhibit an asymmetry in the emission rate with a preference for either one or the other side. However, this asymmetry is not completely fixed during the decay, and some of the trajectories change their preferred direction of emission during the decay process. As shown in Fig. 17(b), the relative strength of the asymmetry peaks at $t_{\star}$ and increases with disorder strength.
Appendix J Boundary Effects
@@
Figure S3: (a) Probability distribution of relative propagating phases $\Delta\xi=(\xi_{i}-\xi_{j})\bmod 2\pi$, which is a wrapped triangular distribution. (b) Probability distribution of relative spin angles $\Delta\phi=(\phi_{i}-\phi_{j})\bmod 2\pi$, for $N=400$ and $10^{3}$ trajectories. When $\Theta=0$, the system is permutation-invariant in QSDMF trajectories, and all spins align along the same direction, resulting in $p(\Delta\phi)=\delta(0)$ as expected. As $\Theta$ increases, the relative propagating phase $\Delta\xi$ becomes more uniformly distributed in $(-\pi,\pi)$. Accordingly, $p(\Delta\phi)$ also becomes more evenly distributed, meaning the global spin alignment of the atoms weakens. Instead, a new spin ordering emerges, with spins aligning according to their position-dependent phases as revealed by the joint probability distribution $p(\Delta\xi,\Delta\phi)$ shown in Sec. V.2. Additionally, we can observe that, under strong disorder, there is a weak tendency for alignment and anti-alignment in the distribution $p(\Delta\phi)$. This can be attributed to the fact that the joint distribution $p(\Delta\xi,\Delta\phi)$, shown in Fig. 8, has less spreading around $\Delta\xi\approx 0$ and $\pi$. For similar reasons as in Appendix H, the ordering around $\Delta\phi\approx\Delta\xi\approx\pm\pi/2$ is weaker than ordering in other directions.
@@
L Reduction of the SDE for Efficient Treatment of Cavity Dynamics
@@
M Irrelevance of Detuning in DTWA
@@
N Comment on Experimental Relevance
Here, using the parameters of a few recent experiments, we discuss possible experimental realizations and tests of our theoretical predictions. Some experiments already exhibit strong spatial disorder. Experiments with weak spatial disorder can be used to study our predictions for regular arrays.
  • In Refs. [37, 40], $\sim 1000$ cesium atoms are chirally coupled to a 1D nanophotonic waveguide. Since the atoms are probabilistically loaded into the trapping potential and the trapping lattice spacing is incommensurate with the resonant wavelength, the resulting atomic separations give random propagation phases that are roughly uniformly distributed in $(0,2\pi)$ [86]. Therefore, this setup corresponds to the strongly disordered case with $\Theta\approx 2\pi$ in our study. Although random propagation phases can be absorbed into a redefinition of the Pauli operators for a chiral waveguide and thus do not affect the dynamics, coupling to a bidirectional waveguide can be realized with nonchiral waveguide or different atomic transitions and used to directly probe our theoretical predictions.
  • In the microwave regime, Ref. [26] realizes 8 superconducting qubits coupled to a 1D waveguide. The qubits are densely spaced, with separations much smaller than the resonant wavelength and a spatial disorder strength $\Theta\ll 1$. Therefore, spatial disorder can be neglected in this setup, while frequency disorder may play a more important role. When scaled to more qubits, this setup can be used directly to probe our predictions with a predefined disorder pattern or ordering behavior of regular lattices discussed in Appendix H.
  • Ref. [36] studies 3 quantum dots coupled to a photonic crystal waveguide. The positional uncertainty in this setup is $\sim\lambda/10$, i.e. $\Theta\sim\pi/5$, which lies in the weakly disordered regime for this small number of emitters.
  • As an outlook for future study, we mention collective emission into 3D free space. Recently, the experiment in Ref. [33] studied a 3D pencil-shaped cloud of $\sim 2000$ laser-cooled 87Rb atoms under external driving. The size of the atomic cloud is $\sim 0.5\lambda$ and $\sim 20\lambda$ along the two axes. This gives large disorder strengths $\Theta\sim\pi$ and $40\pi$ in these two directions. Our theory can potentially be extended to this quasi-1D case and probed in such experiments.